Annapolis Company has two service departments (Computer Operations & Maintenance Services). Annapolis has two production departments (Mixing Department & Packaging Department.) Annapolis uses a direct allocation method where service departments are allocated only to production departments. All allocations are based on total employees. Computer Operations has costs of $195,000 and Maintenance Services has costs of $150,000 before any allocations. What amount of Maintenance Services total cost is allocated to the Packaging Department? (round to closest whole dollar) Employees are:

            Computer Operations 2

            Maintenance Services 2

Mixing Department 6

Packaging Department 5
Answer #1

Total allocable employees = 6+5 = 11

Maintenance services cost allocated to Packaging department = 150000*5/11 = 68182